# Anton D. Utyushev,a Vadim I. Zakomirnyi,b Alexey A. Shcherbakov,a Ilia L. Rasskazovc and Alexander Morozd

### v1

@@ 3.1 The dipole MLWA vs perturbation theory
Fig.  4: Extinction spectra, $Q_{\rm ext}$, for Al, Ag, Au and Mg spherical nanoparticles with $a=30$ nm embedded in a host absorption medium with $n_{2}^{\prime}=1.33$ (left column) and gain host medium (right column). The left (right) column is for $\bar{\kappa}=0.01$ ($\bar{\kappa}=-0.01$). $Q_{\rm ext}$ are shown as calculated by eqn (1) of the Mie theory (solid red line), within dipole MLWA of eqn (3) (blue dot-dashed line), the Mie theory approximation with the nonabsorbing host (green dashed line), and by the $\bar{\kappa}$-PT given by eqn (9) (dotted orange line). The refractive indices of Ag, Al, Au were taken from ref. 35 and those of Mg from ref. 36.
@@
  • the dipole MLWA (with the sole term $T_{E1}$ in eqn (1) given by eqn (3))
  • the Mie theory approximation with nonabsorbing host (determined by eqn (1) with $n_{2}^{\prime}=0$)
  • the first-order $\bar{\kappa}$-PT given by eqn (9)
@@
in water-like host ($n_{2}^{\prime}=1.33$) for different host absorption characterized by different values of $\bar{\kappa}$ on the example of Ag spheres with radii between $10$ and $50$ nm. Fig. 4 compares results for spheres made from most common plasmonic materials Al, Ag, Au, Mg with radius $a=30$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=0.01$ and $-0.01$ (similar results for glass-like host ($n_{2}^{\prime}=1.5$) are presented in Fig. S2I of ESI). As demonstrated in those figures, the dipole MLWA can be very precise. In the case of Ag and Au nanoparticles, the dipole MLWA essentially overlies with the exact Mie theory results for $a\lesssim 25$ nm. The agreement, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm (Fig. S1III of ESI), and can be used, at least qualitatively, up to $a\sim 70$ nm (Fig. S1IV of ESI). In the case of Al and Mg nanoparticles, a slight deviation from Mie theory results becomes visible by the naked eye already for $a\gtrsim 25$ nm (cf. Figs. S1I and S1II of ESI). The agreement continues to be acceptable up to $a\sim 40$ nm and can be used at least qualitatively up to $a\sim 50$ nm (Fig. S1III of ESI).
@@
A first general observation is that with increasing sphere radius the performance of the dipole MLWA worsens. This is not surprising, as the MLWA is by definition is a long wavelength approximation. The same applies to the $\bar{\kappa}$-PT. The $\bar{\kappa}$-PT overlies with the dipole MLWA for all $a\leq 50$ nm, provided that $\bar{\kappa}\leq 0.01$. The latter justifies a posteriori that our $\bar{\kappa}$-PT is correct. Another interesting tendency revealed by Fig. 3 (Figs. S1I and S1II of ESI) is that the maximal radius $a$ for which the dipole MLWA remains accurate increases with increasing $\bar{\kappa}$.
@@
Fig.  5: Size dependence of $\bar{Q}_{\bar{\kappa}}$, the ratio of the PT correction to the leading term on the rhs of eqn (9), for Ag sphere at the (size-dependent) LSPR wavelength for different $\bar{\kappa}$: (a) $0.001$, (b) $0.005$, (c) $0.01$ and (d) $0.1$.
@@
Fig.  6: Extinction spectra, $Q_{\rm ext}$, for spherical Ag nanoparticles with (a) $a=50$ nm, (b) $a=60$ nm, (c) $a=70$ nm, and (d) $a=80$ nm embedded in Poly(3-hexylthiophene) (P3HT) host medium. Spectra are shown as calculated with the exact Mie theory (solid red line), the Mie theory for non-absorbing media ($n_{2}^{\prime\prime}=0$) (green dashed line), the dipole MLWA (blue dot-dashed line), and the $\bar{\kappa}$-PT (dotted orange line).
@@
Importantly, the results of $\bar{\kappa}$-PT and the dipole MLWA overlie with the exact Mie theory results for $a\lesssim 25$ nm, which supports earlier observation on the range of validity of the MLWA in an absorbing host by Khlebtsov 17. Whereas the dipole MLWA continues to overlie with the exact Mie theory results for $a\lesssim 25$ nm irrespective of $\bar{\kappa}$, the results of the $\bar{\kappa}$-PT begin to deviate from those of the dipole MLWA for $\bar{\kappa}\gtrsim 0.05$. This is to be expected, because the $\bar{\kappa}$-PT is a first order perturbation theory of the dipole MLWA in $\bar{\kappa}$.
@@
Surprisingly, the smaller the particle, the greater the deviation of the $\bar{\kappa}$-PT relative to the dipole MLWA for $\bar{\kappa}\gtrsim 0.05$ (see Fig. S1 and S2 of ESI). The origin of this behaviour is that $D(\varepsilon_{t})$ is typically small in a proximity of a LSPR (cf. the generalized Fröhlich condition eqn (5)). Whereas the first term in eqn (9) is of the order $1/D(\varepsilon_{t})$, the term proportional to $\bar{\kappa}$ is of the order $1/D^{2}(\varepsilon_{t})$. However, when the size parameter $x$ increases, $|D(\varepsilon_{t})|$ at a LSPR decreases (due to a larger separation from the complex zero), and, as illustrated in Fig. 5, $\bar{Q}_{\bar{\kappa}}$, defined as the ratio of the PT correction to the leading term on the rhs of eqn (9), gradually decreases, whereby the first-order $\bar{\kappa}$-PT begins to approximate the MLWA. Note also how the ratio $\bar{Q}_{\bar{\kappa}}$ increases with $\bar{\kappa}$, which, as expected, explains why the $\bar{\kappa}$-PT begins to deviate from the MLWA with increasing $\bar{\kappa}$.
@@ 3.2 Mie theory approximation with nonabsorbing host
Similar behaviour is observed for the Mie theory approximation with nonabsorbing host ($n_{2}^{\prime\prime}=0$) that begins to deviate from the exact Mie theory results, but somewhat earlier, beginning with $\bar{\kappa}=0.01$. Again the smaller the particle, the greater the deviation. Nevertheless, the Mie theory approximation with nonabsorbing host becomes the best approximation with increasing particle radius (e.g. $a\gtrsim 40$ nm and $\bar{\kappa}\leq 0.01$). For example, the Mie theory in nonabsorbing host begins to overlie with the Mie theory in an absorbing host for $a\gtrsim 50$ nm and $\bar{\kappa}=0.01$ (see Fig. S1 of ESI). A threshold radius for which it happens increases with $\bar{\kappa}$.
Why the Mie theory approximation with nonabsorbing host becomes the best approximation for sufficiently large $a$ can be explained by increasing relevance of higher-order multipole contributions with increasing particle size. This is understandable, because with increasing $a$ higher order multipoles, which are obviously absent in any dipole approximation, become more and more relevant. This is demonstrated also in Fig. 6 showing the effect of increasing the radius of spherical Ag nanoparticles in Poly(3-hexylthiophene) (P3HT) host medium, which was used in a number of recent studies 15, 37, 17, 19. Unlike inherent dipole approximations, the Mie theory approximation with nonabsorbing host captures reasonably well both the dipole and quadrupole peaks. Obviously, higher-order MLWA of ref. 34 could have captured the quadrupole peak, but this goes beyond the scope of present study.
@@ 3.3 Isolating the host dissipative effects on the extinction cross sections
The dissipative effects of the host on the extinction cross sections, $Q_{\rm ext}$, can obviously be isolated by subtracting from the exact value $Q_{\rm ext}=Q_{{\rm ext};\varepsilon}$ the value of $Q_{{\rm ext};\varepsilon_{t}}$ obtained by the Mie theory for the non-absorbing host characterised by $n_{2}^{\prime\prime}=0$ (i.e. $\varepsilon\to\varepsilon_{t}$). In what follows, we denote the difference by $\eta_{\rm eff}:=Q_{{\rm ext};\varepsilon}-Q_{{\rm ext};\varepsilon_{t}}$. Thus $\eta_{\rm eff}$ quantifies the contribution of the host absorption or gain in the resulting $Q_{\rm ext}$. To our satisfaction, it turns out that, in a suitable parameter range, the analytic $\bar{\kappa}$-PT can reliably capture the effect of the host absorption on the extinction efficiency of a plasmonic nanosphere as demonstrated in Fig. 7. Not surprisingly, the agreement can also be reached in the case of gain media, which opens door for analysis of promising applications involving active media, see Fig. 4 (Fig. S5 of ESI) such as spasers 38, 39, 40. Therefore, within the range of its validity, the first order $\bar{\kappa}$-PT allows one to both intuitively and analytically understand the mechanisms of host dissipation or gain on the extinction efficiency of a plasmonic nanosphere.
Fig.  7: The effect $\eta_{\rm eff}$ of the host absorption (left column) and the gain host medium (right column) on $Q_{\rm ext}$ in the Mie theory (solid red line) and the $\bar{\kappa}$-PT (dashed dark green line) for Al, Ag, Au and Mg materials with different radii $a$ embedded in the host medium with $n_{2}^{\prime}=1.33$ and $\bar{\kappa}=0.001$ (left column) and $\bar{\kappa}=-0.001$ (right column).
@@ 4 Discussion
Rather exhaustive MLWA analysis in the absorbing case has been recently provided by Khlebtsov 17 by employing slightly different MLWA. The study of ref. 17 arrived at similar conclusions that MLWA is very precise for $a\lesssim 25$ nm. Our study complements ref. 17 with (i) a first order analytic perturbation theory around a nonabsorbing host in a normalized extinction coefficient $\bar{\kappa}$ ($\bar{\kappa}$-PT) and (ii) an investigation of gain media. In addition, we have examined the behaviour of Al and Mg nanoparticles.
@@ 5 Conclusions
Our previously developed modified dipole long-wave approximation (MLWA) was shown to essentially overly with the exact Mie theory results for $a\lesssim 25$ nm ($a\lesssim 20$ nm) in the case of Ag and Au (Al and Mg) nanoparticles. The agreement for Au and Ag (Al and Mg) nanoparticles, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm ($a\sim 40$ nm), and can be used, at least qualitatively, up to $a\sim 70$ nm ($a\sim 50$ nm). We developed within the dipole MLWA a first order analytic perturbation theory (PT) around a nonabsorbing host in a normalized extinction coefficient $\bar{\kappa}$ and investigated its properties. It was shown that, in a suitable parameter range, the $\bar{\kappa}$-PT can reliably capture the effect of host absorption or gain on the extinction efficiency of a plasmonic nanosphere. Considering growing interest in light-matter interactions, we expect that our results will help in designing optimal systems comprising plasmonic nanoparticles embedded in suitable dissipative or gain media for various applications.
Author Contributions
Conflicts of interest
@@
Acknowledgements
Electronic Supplementary Information
The supplementary information includes Sec. S1 “Supplementary Figures”, where additional visual data are provided on apparent extinction efficiencies for further particle radii, host $n_{2}^{\prime}$, and normalized extinction coefficient $\bar{\kappa}$. In Sec. S2 “Complex Pole of the Dipole Contribution within the Dipole MLWA”, the predictive power of the complex pole knowledge on the LSPR position is investigated. Sec. S3 “Proof of that $\varepsilon_{E1}$ is the Complex Zero of the Denominator of $a_{1}$” presents mathematical proof of what is in its title. In Sec. S4 “The Imaginary Part of $\varepsilon_{E1}$ is Negative” proves the stated statement for any real $x$. In Sec. S5 “The Derivative $da_{1}/d\bar{\kappa}$” explicit calculations leading to eqn (7) are presented. Sec. S6 “A Previous Attempt to Formulate a Perturbation Theory (PT) in Terms of $\bar{\kappa}$” analyses an earlier failed theoretical attempt. Sec. S7 “Apparent Extinction Cross Section” reviews the absorption and scattering cross-sections within the MLWA. Finally, Sec. S8 “Conventional Scattering Theory” reviews some related aspects of conventional scattering theory.
Notes and references
  • Newton 1982 R. G. Newton, Scattering Theory of Waves and Particles, Springer Berlin Heidelberg, Berlin, Heidelberg, 1982.
  • Bohren and Huffman 1998 C. F. Bohren and D. R. Huffman, Absorption and Scattering of Light by Small Particles, Wiley-VCH Verlag GmbH & Co. KGaA, 1998.
  • Fuchs and Kliewer 1968 R. Fuchs and K. L. Kliewer, J. Opt. Soc. Am., 1968, 58, 319–330.
  • Mundy et al. 1974 W. C. Mundy, J. A. Roux and A. M. Smith, J. Opt. Soc. Am., 1974, 64, 1593–1597.
  • Chýlek and Pinnick 1979 P. Chýlek and R. G. Pinnick, Appl. Opt., 1979, 18, 1123–1124.
  • Bohren and Gilra 1979 C. F. Bohren and D. P. Gilra, J. Colloid Interface Sci., 1979, 72, 215–221.
  • Lebedev et al. 1999 A. N. Lebedev, M. Gartz, U. Kreibig and O. Stenzel, Eur. Phys. J. D., 1999, 6, 365–373.
  • Sudiarta and Chylek 2001 I. W. Sudiarta and P. Chylek, J. Opt. Soc. Am. A, 2001, 18, 1275–1278.
  • Yang et al. 2002 P. Yang, B.-C. Gao, W. J. Wiscombe, M. I. Mishchenko, S. E. Platnick, H.-L. Huang, B. A. Baum, Y. X. Hu, D. M. Winker, S.-C. Tsay and S. K. Park, Appl. Opt., 2002, 41, 2740–2759.
  • Fu and Sun 2006 Q. Fu and W. Sun, J. Quant. Spectrosc. Radiat. Transf., 2006, 100, 137–142.
  • Mishchenko 2007 M. I. Mishchenko, Opt. Express, 2007, 15, 13188–13201.
  • Mishchenko et al. 2017 M. I. Mishchenko, G. Videen and P. Yang, Opt. Lett., 2017, 42, 4873–4876.
  • Mishchenko and Yang 2018 M. I. Mishchenko and P. Yang, J. Quant. Spectrosc. Radiat. Transf., 2018, 205, 241–252.
  • Mishchenko and Dlugach 2019 M. I. Mishchenko and J. M. Dlugach, Appl. Opt., 2019, 58, 4871–4877.
  • Peck et al. 2019 R. L. Peck, A. G. Brolo and R. Gordon, J. Opt. Soc. Am. B, 2019, 36, F117–F122.
  • Mishchenko et al. 2019 M. I. Mishchenko, M. A. Yurkin and B. Cairns, OSA Continuum, 2019, 2, 2362–2368.
  • Khlebtsov 2021 N. G. Khlebtsov, Phys. Chem. Chem. Phys., 2021, 23, 23141–23157.
  • Dong et al. 2021 J. Dong, W. Zhang and L. Liu, Opt. Express, 2021, 29, 7690–7705.
  • Zhang et al. 2022 S. Zhang, J. Dong, W. Zhang, Minggang Luo and L. Liu, Opt. Lett., 2022, 47, 5577–5580.
  • Mishchenko and Dlugach 2019 M. I. Mishchenko and J. M. Dlugach, OSA Continuum, 2019, 2, 3415–3421.
  • Videen and Sun 2003 G. Videen and W. Sun, Appl. Opt., 2003, 42, 6724–6727.
  • Meier and Wokaun 1983 M. Meier and A. Wokaun, Opt. Lett., 1983, 8, 581–583.
  • Zeman and Schatz 1984 E. J. Zeman and G. C. Schatz, Dynamics on Surfaces, Dordrecht, 1984, pp. 413–424.
  • Zeman and Schatz 1987 E. J. Zeman and G. C. Schatz, J. Phys. Chem., 1987, 91, 634–643.
  • Kelly et al. 2003 K. L. Kelly, E. Coronado, L. L. Zhao and G. C. Schatz, J. Phys. Chem. B, 2003, 107, 668–677.
  • Kuwata et al. 2003 H. Kuwata, H. Tamaru, K. Esumi and K. Miyano, Appl. Phys. Lett., 2003, 83, 4625–4627.
  • Moroz 2009 A. Moroz, J. Opt. Soc. Am. B, 2009, 26, 517–527.
  • Zorić et al. 2011 I. Zorić, M. Zäch, B. Kasemo and C. Langhammer, ACS Nano, 2011, 5, 2535–2546.
  • Massa et al. 2013 E. Massa, S. A. Maier and V. Giannini, New J. Phys., 2013, 15, 063013.
  • Le Ru et al. 2013 E. C. Le Ru, W. R. C. Somerville and B. Auguié, Phys. Rev. A, 2013, 87, 012504.
  • Schebarchov et al. 2013 D. Schebarchov, B. Auguié and E. C. Le Ru, Phys. Chem. Chem. Phys., 2013, 15, 4233–4242.
  • Januar et al. 2020 M. Januar, B. Liu, J.-C. Cheng, K. Hatanaka, H. Misawa, H.-H. Hsiao and K.-C. Liu, J. Phys. Chem. C, 2020, 124, 3250–3259.
  • Rasskazov et al. 2020 I. L. Rasskazov, P. S. Carney and A. Moroz, Opt. Lett., 2020, 45, 4056–4059.
  • Rasskazov et al. 2021 I. L. Rasskazov, V. I. Zakomirnyi, A. D. Utyushev, P. S. Carney and A. Moroz, J. Phys. Chem. C, 2021, 125, 1963–1971.
  • McPeak et al. 2015 K. M. McPeak, S. V. Jayanti, S. J. P. Kress, S. Meyer, S. Iotti, A. Rossinelli and D. J. Norris, ACS Photonics, 2015, 2, 326–333.
  • Palm et al. 2018 K. J. Palm, J. B. Murray, T. C. Narayan and J. N. Munday, ACS Photonics, 2018, 5, 4677–4686.
  • Peck et al. 2021 R. Peck, A. Khademi, J. Ren, S. Hughes, A. G. Brolo and R. Gordon, Phys. Rev. Research, 2021, 3, 013014.
  • Bergman and Stockman 2003 D. J. Bergman and M. I. Stockman, Phys. Rev. Lett., 2003, 90, 027402.
  • Lawandy 2004 N. M. Lawandy, Appl. Phys. Lett., 2004, 85, 5040–5042.
  • Veltri et al. 2016 A. Veltri, A. Chipouline and A. Aradian, Sci. Rep., 2016, 6, 33018.

### v2

@@ 3.1 The dipole MLWA vs perturbation theory
Figure 4: Extinction spectra, $Q_{\rm ext}$, for Al, Ag, Au and Mg spherical nanoparticles with $a=25$ nm embedded in a host absorption medium with $n_{2}^{\prime}=1.33$ (left two columns) and gain host medium (right two columns) for $\bar{\kappa}=0.01,0.1$ ($\bar{\kappa}=-0.01,-0.1$). $Q_{\rm ext}$ are shown as calculated by Eq. (1) of the Mie theory (solid red line), within dipole MLWA of Eq. (3) (blue dot-dashed line), the Mie theory approximation with the nonabsorbing host (green dashed line), and by the $\bar{\kappa}$-PT given by Eq. (9) (dotted orange line). The refractive indices of Ag, Al, Au were taken from ref. 35 and those of Mg from ref. 36. Whereas absorption decreases the height of a LSPR and broadens its width, the gain does just the opposite. Note enlarged scale on the vertical axis in the last column for $\bar{\kappa}=-0.1$.
@@
  • the dipole MLWA (with the sole term $T_{E1}$ in Eq. (1) given by Eq. (3))
  • the Mie theory approximation with nonabsorbing host (determined by Eq. (1) with $n_{2}^{\prime\prime}=0$)
  • the first-order $\bar{\kappa}$-PT given by Eq. (9)
@@
in water-like host ($n_{2}^{\prime}=1.33$) for different host absorption characterized by different values of $\bar{\kappa}$ on the example of Ag spheres with radii between $10$ and $50$ nm. Fig. 4 compares results for spheres made from most common plasmonic materials Al, Ag, Au, Mg with radius $a=25$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=\pm 0.01$ and ±$\pm 0.1$ (similar results for glass-like host ($n_{2}^{\prime}=1.5$) are presented in Fig. S2 in Supplement 1). For different metals of plasmonic particles tabulated data of refractive indices have been used as follows: Al [35], Ag [35], Au [35], Mg [36]. As demonstrated in those figures, the dipole MLWA can be very precise. In the case of Ag and Au nanoparticles, the dipole MLWA essentially overlies with the exact Mie theory results for $a\lesssim 25$ nm. The agreement, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm (Fig. S1III in Supplement 1), and can be used, at least qualitatively, up to $a\sim 70$ nm (Fig. S1IV in Supplement 1). In the case of Al and Mg nanoparticles, a slight deviation from Mie theory results becomes visible by the naked eye already for $a\gtrsim 25$ nm (cf. Figs. S1I and S1II in Supplement 1). The agreement continues to be acceptable up to $a\sim 40$ nm and can be used at least qualitatively up to $a\sim 50$ nm (Fig. S1III in Supplement 1).
@@
A first general observation is that with increasing sphere radius the performance of the dipole MLWA worsens. This is not surprising, as the MLWA is by definition is a long wavelength approximation. The same applies to the $\bar{\kappa}$-PT. The $\bar{\kappa}$-PT overlies with the dipole MLWA for all $a\leq 50$ nm, provided that $\bar{\kappa}\leq 0.01$. The latter justifies a posteriori that our $\bar{\kappa}$-PT is correct. Another interesting tendency revealed by Fig. 3 (Figs. S1I and S1II in Supplement 1) is that the maximal radius $a$ for which the dipole MLWA remains accurate increases with increasing $\bar{\kappa}$. This observation can be explained by the fact that for a given particle radius an increasing host absorption suppresses the contribution of higher-order multipoles that would normally arise at larger particle sizes. This allows the dipole MLWA to remain valid over a wider range of particle sizes, as the multipole effects are suppressed in absorbing media.
@@
Figure 5: Size dependence of $\bar{Q}_{\bar{\kappa}}$, the ratio of the PT correction to the leading term on the rhs of Eq. (9), for Ag sphere at the (size-dependent) LSPR wavelength for different $\bar{\kappa}$: (a) $0.001$, (b) $0.005$, (c) $0.01$ and (d) $0.1$.
@@
Figure 6: Extinction spectra, $Q_{\rm ext}$, for spherical Ag nanoparticles with (a) $a=50$ nm, (b) $a=60$ nm, (c) $a=70$ nm, and (d) $a=80$ nm embedded in Poly(3-hexylthiophene) (P3HT) host medium. Spectra are shown as calculated with the exact Mie theory (solid red line), the Mie theory for non-absorbing media ($n_{2}^{\prime\prime}=0$) (green dashed line), the dipole MLWA (blue dot-dashed line), and the $\bar{\kappa}$-PT (dotted orange line).
@@
Importantly, the results of $\bar{\kappa}$-PT and the dipole MLWA overlie with the exact Mie theory results for $a\lesssim 25$ nm, which supports earlier observation on the range of validity of the MLWA in an absorbing host by Khlebtsov [17]. Whereas the dipole MLWA continues to overlie with the exact Mie theory results for $a\lesssim 25$ nm irrespective of $\bar{\kappa}$, the results of the $\bar{\kappa}$-PT begin to deviate from those of the dipole MLWA for $\bar{\kappa}\gtrsim 0.05$. An extreme case of such a deviation is provided for Ag sphere with $a=25$ nm and $\bar{\kappa}=\pm 0.1$ shown in Fig. 4(f),(h). Whereas the dipole MLWA nearly overlies the exact result showing a positive (negative) extinction for $\bar{\kappa}=0.1$ ($\bar{\kappa}=-0.1$), the $\bar{\kappa}$-PT in the respective cases indicates a negative (positive) extinction. This is to be expected, because the $\bar{\kappa}$-PT is a first order perturbation theory of the dipole MLWA in $\bar{\kappa}$ and is expected to eventually break down above a certain threshold value of $|\bar{\kappa}|$. Negative extinction shown in Fig. 4(h) is an indication of that particle losses have been more than compensated by the gain medium - a precursor of lasing action [37, 38, 39].
@@
Surprisingly, the smaller the particle, the greater the deviation of the $\bar{\kappa}$-PT relative to the dipole MLWA for $\bar{\kappa}\gtrsim 0.05$ (see Fig. S1 and S2 in Supplement 1). The origin of this behaviour is that $D(\varepsilon_{t})$ is typically small in a proximity of a LSPR (cf. the generalized Fröhlich condition Eq. (5)). Whereas the first term in Eq. (9) is of the order $1/D(\varepsilon_{t})$, the term proportional to $\bar{\kappa}$ is of the order $1/D^{2}(\varepsilon_{t})$. However, when the size parameter $x$ increases, $|D(\varepsilon_{t})|$ at a LSPR decreases (due to a larger separation from the complex zero), and, as illustrated in Fig. 5, $\bar{Q}_{\bar{\kappa}}$, defined as the ratio of the PT correction to the leading term on the rhs of Eq. (9), gradually decreases, whereby the first-order $\bar{\kappa}$-PT begins to approximate the MLWA. Note also how the ratio $\bar{Q}_{\bar{\kappa}}$ increases with $\bar{\kappa}$, which, as expected, explains why the $\bar{\kappa}$-PT begins to deviate from the MLWA with increasing $\bar{\kappa}$.
@@ 3.2 Mie theory approximation with nonabsorbing host
Similar behaviour is observed for the Mie theory approximation with nonabsorbing host ($n_{2}^{\prime\prime}=0$) that begins to deviate from the exact Mie theory results, but somewhat earlier, beginning with $\bar{\kappa}=0.01$. Again the smaller the particle, the greater the deviation. Nevertheless, in a relatively weakly absorbing host the Mie theory approximation with nonabsorbing host becomes the best approximation. For example, the Mie theory in nonabsorbing host begins to overlie with the Mie theory in an absorbing host for $a\gtrsim 50$ nm and $\bar{\kappa}=0.01$ (see Fig. S1 in Supplement 1). A threshold radius for which it happens increases with $\bar{\kappa}$.
Why the Mie theory approximation with nonabsorbing host becomes the best approximation for sufficiently large $a$ can be explained by increasing relevance of higher-order multipole contributions with increasing particle size. This is understandable, because with increasing $a$ higher order multipoles, which are obviously absent in any dipole approximation, become more and more relevant. This is demonstrated also in Fig. 6 showing the effect of increasing the radius of spherical Ag nanoparticles in Poly(3-hexylthiophene) (P3HT) host medium, which was used in a number of recent studies [15, 40, 17, 19]. Unlike inherent dipole approximations, the Mie theory approximation with nonabsorbing host captures reasonably well both the dipole and quadrupole peaks. Obviously, higher-order MLWA of ref. 34 could have captured the quadrupole peak, but this goes beyond the scope of present study.
@@ 3.3 Isolating the host dissipative effects on the extinction cross sections
The dissipative effects of the host on the extinction cross sections, $Q_{\rm ext}$, can obviously be isolated by subtracting from the exact value $Q_{\rm ext}=Q_{{\rm ext};\varepsilon}$ the value of $Q_{{\rm ext};\varepsilon_{t}}$ obtained by the Mie theory for the non-absorbing host characterised by $n_{2}^{\prime\prime}=0$ (i.e. $\varepsilon\to\varepsilon_{t}$). In what follows, we denote the difference by $\eta_{\rm eff}:=Q_{{\rm ext};\varepsilon}-Q_{{\rm ext};\varepsilon_{t}}$. Thus $\eta_{\rm eff}$ quantifies the contribution of the host absorption or gain in the resulting $Q_{\rm ext}$. To our satisfaction, it turns out that, in a suitable parameter range, the analytic $\bar{\kappa}$-PT can reliably capture the effect of the host absorption on the extinction efficiency of a plasmonic nanosphere as demonstrated in Fig. 7. Not surprisingly, the agreement can also be reached in the case of gain media, which opens door for analysis of promising applications involving active media, see Fig. 4 (Fig. S5 in Supplement 1) such as spasers [37, 38, 39]. Therefore, within the range of its validity, the first order $\bar{\kappa}$-PT allows one to both intuitively and analytically understand the mechanisms of host dissipation or gain on the extinction efficiency of a plasmonic nanosphere.
Figure 7: The effect $\eta_{\rm eff}$ of the host absorption (left column) and the gain host medium (right column) on $Q_{\rm ext}$ in the Mie theory (solid red line) and the $\bar{\kappa}$-PT (dashed dark green line) for Al, Ag, Au and Mg materials with different radii $a$ embedded in the host medium with $n_{2}^{\prime}=1.33$ and $\bar{\kappa}=0.001$ (left column) and $\bar{\kappa}=-0.001$ (right column).
@@ 4 Discussion
First-principles far-field computations based on the general Lorenz–Mie theory showed that increasing absorption in the host medium broadens and suppresses plasmon resonances in the apparent extinction [14]. Such a broadening and suppression of plasmon resonances is also present in all approximation shown in Figs. 3 and  4. The effect of absorption in the host thus goes in the same direction as increasing absorption within the particle. Noteworthy, increasing gain in the host medium narrows plasmon resonances and increases their amplitude in the apparent extinction as demonstrated in Fig. 4 (cf. also Fig. S3). This goes along physical intuition. Contrary to the apparent extinction, in the case of the inherent cross-sections a surrounding lossy medium was shown to narrow the plasmonic resonances, as well as increase their amplitude dramatically [15, 17].
The dipole MLWA can, in principle, be used up to the laser threshold associated with the so-called spectral singularity of non-Hermitian models. Indeed, as it has been verified both experimentally and theoretically more than 15 years ago, a description of gain media consisting in changing the sign of extinction is a valid approximation up to the laser threshold [41]. This could be particularly important for applications in spasers and nano-lasers.
For the sake of completeness, rather exhaustive MLWA analysis in the absorbing case has been recently provided by Khlebtsov [17] by employing slightly different MLWA. The study of ref. 17 arrived at similar conclusions that MLWA is very precise for $a\lesssim 25$ nm. Our study complements ref. 17 with (i) a first order analytic perturbation theory around a nonabsorbing host in a normalized extinction coefficient $\bar{\kappa}$ ($\bar{\kappa}$-PT) and (ii) an investigation of gain media. In addition, we have examined the behaviour of Al and Mg nanoparticles.
@@ 5 Conclusions
Our previously developed modified dipole long-wave approximation (MLWA) was shown to essentially overly with the exact Mie theory results for $a\lesssim 25$ nm ($a\lesssim 20$ nm) in the case of Ag and Au (Al and Mg) nanoparticles. The agreement for Au and Ag (Al and Mg) nanoparticles, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm ($a\sim 40$ nm), and can be used, at least qualitatively, up to $a\sim 70$ nm ($a\sim 50$ nm). Expanding the discussion to the case of larger nanoparticles with $a>70$ nm almost certainly requires consideration of MLWA for higher order multipoles [31, 34]. We developed within the dipole MLWA a first order analytic perturbation theory (PT) around a nonabsorbing host in a normalized extinction coefficient $\bar{\kappa}$ and investigated its properties. It was shown that, in a suitable parameter range, the $\bar{\kappa}$-PT can reliably isolate and capture the effect of host absorption or host gain on the overall extinction efficiency of spherical plasmonic nanoparticles. Considering growing interest in light-matter interactions, we expect that our results will help in designing optimal systems comprising plasmonic nanoparticles embedded in suitable dissipative or gain media for various applications, such as photothermal therapy [42, 43], spasers and other active medium devices [37, 38, 39].
6 Funding
The analytical study of absorbing media is funded by Russian Science Foundation grant № 23-19-00511.
7 Declaration of competing interest
@@
8 Data availability
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
9 Supplemental document
See Supplement materials for supporting content.